# Empowering AlphaFold2 for protein conformation selective drug discovery with AlphaFold2-RAVE

### v2

@@ 2.5 Discussion
With the rapid expansion of chemical space, virtual screening on libraries containing billions of diverse molecules becomes enticing for novel drug discovery 41. Therefore, the enrichment of candidate holo-like structures emerges as a necessary step, offering significant benefits in terms of computational efficiency and feasibility. As summarized in Table 2.5, unlike the non-dockable AF2 structures for type II inhibitors, the diverse rMSA AF2 ensemble (referred to as rAF2 in Table 2.5 for brevity) shows potential in generating holo-like structures within a large set of decoys. However, it’s only upon AF2RAVE ranking and selection that the ratio of holo-like structures in selected structure set increases to a plausible value of 50%, facilitating further virtual screening on computational models of protein pockets.
@@ 3 Conclusion
For the development of novel drugs targeting general proteins, all metastable states identified by AF2RAVE should be explored in subsequent docking experiments. Integration of algorithms capable of predicting ligand binding sites on protein surfaces, such as the Graph Attention Site Prediction (GrASP) model42, is then essential before utilizing AF2RAVE-selected structures in docking, thus expanding the workflow to AF2RAVE-GrASP-Glide. Additionally, the inclusion of free energy perturbation (FEP) calculations for front-runner ligands to evaluate the actual ligand binding affinity can further enhance this workflow.
@@
We conclude this manuscript by making a final comment on the horizons opened by Generative AI methods, including those involving diffusion models10, 43 and any such future frameworks. These approaches make it possible to easily hypothesize regions of the conformational space underlying arbitrarily complex molecules of life, that then could serve as a starting point to launch more careful investigations. However, these predictions without careful in situ or a posteriori testing through advanced simulations or experiments, are only predictions and could just be hallucinations. Having the capability to quickly generate numerous - thousands or more - such structural hypotheses is what made AlphaFold2 so crucial to this current work. We believe that a deep integration of the hypothesis creation possibilities of Generative AI with careful Molecular Dynamics,44, 45 experiments or other forms of rapid testing is one way these methods will truly facilitate new and reliable scientific discoveries. {acknowledgement}
@@
Research in this publication was supported by the National Institute of General Medical Sciences of the National Institutes of Health under Award Number R35GM142719. The content is solely the responsibility of the authors and does not represent the official views of the National Institutes of Health. We thank UMD HPC’s Zaratan and NSF ACCESS (project CHE180027P) for computational resources. A.A. was supported by NCI-UMD Partnership for Integrative Cancer Research. P.T. is an investigator at the University of Maryland-Institute for Health Computing, which is supported by funding from Montgomery County, Maryland and The University of Maryland Strategic Partnership: MPowering the State, a formal collaboration between the University of Maryland, College Park and the University of Maryland, Baltimore. We thank UMD HPC’s Zaratan and NSF ACCESS (project CHE180027P) for computational resources. We thank Bodhi Vani, Dedi Wang and Zachary Smith for helpful discussions.
@@ 3.1 Notes
Detailed description of methods used and further analysis of systems and sampling can be found in the supplement.

### v3

@@ 2.5 Discussion
With the rapid expansion of chemical space, virtual screening on libraries containing billions of diverse molecules becomes enticing for novel drug discovery 44. Therefore, the enrichment of candidate holo-like structures emerges as a necessary step, offering significant benefits in terms of computational efficiency and feasibility. As summarized in Table 2.5, unlike the non-dockable AF2 structures for type II inhibitors, the diverse rMSA AF2 ensemble (referred to as rAF2 in Table 2.5 for brevity) shows potential in generating holo-like structures within a large set of decoys. However, it’s only upon AF2RAVE ranking and selection that the ratio of holo-like structures in selected structure set increases to a plausible value of 50%, facilitating further virtual screening on computational models of protein pockets.
@@ 3 Conclusion
This demonstration of AF2RAVE-Glide on kinase inhibitors shows its promising application for discovering drugs targeting general proteins in addition to kinases, such as G-Protein Coupled Receptors (GPCRs), which are the targets for over one-third of Food and Drug Administration (FDA)-approved drugs45. For the design of novel drugs targeting general proteins, developing a protocol that does not require prior information about the system is left for future exploration. Besides, in this study we only investigate the classical DFG-out metastable state for kinases. For a comprehensive protocol, all top-ranked metastable states identified by AF2RAVE should be explored in subsequent docking experiments. Integration of algorithms capable of predicting ligand binding sites on protein surfaces, such as the Graph Attention Site Prediction (GrASP) model46, is then essential before utilizing AF2RAVE-selected structures in docking, thus expanding the workflow to AF2RAVE-GrASP-Glide. Additionally, the inclusion of free energy perturbation (FEP) calculations for front-runner ligands to evaluate the actual ligand binding affinity can further enhance this workflow.
@@
We conclude this manuscript by making a final comment on the horizons opened by Generative AI methods, including those involving diffusion models10, 47 and any such future frameworks. These approaches make it possible to easily hypothesize regions of the conformational space underlying arbitrarily complex molecules of life, that then could serve as a starting point to launch more careful investigations. However, these predictions without careful in situ or a posteriori testing through advanced simulations or experiments, are only predictions and could just be hallucinations. Having the capability to quickly generate numerous - thousands or more - such structural hypotheses is what made AlphaFold2 so crucial to this current work. We believe that a deep integration of the hypothesis creation possibilities of Generative AI with careful Molecular Dynamics,48, 49 experiments or other forms of rapid testing is one way these methods will truly facilitate new and reliable scientific discoveries.
4 Methods
4.1 Summary of systems tested and tools used
In this work we tested AF2RAVE-Glide protocol on active and inactive conformations of DDR1, Abl1 and Src kinases against type I inhibitors VX-680 and Dasatinib, and type II inhibitors Imatinib and Ponatinib. The top-ranked structures obtained from AF2RAVE-Glide were compared against publicly available crystal structures deposited in the Protein Data Bank (PDB), with PDB codes provided in the main text.
To generate diverse structural ensembles for DDR1, Abl1, and Src kinases, we primarily used the reduced MSA AF2 (rMSA AF2) method. For comparison, we also produced AF2-cluster ensembles for these three kinases.
For MD simulations, all the systems in this paper were parameterized with the Amber99SB*-ILDN force field 50, 51 with the TIP3P water model52 and neutralized with Na+ ions or Cl- ions. The simulations are performed at 300 K with the LangevinMiddleIntegrator53 in OpenMM54 with the step size of 2 fs. Particle Mesh Ewald55 is used for calculating electrostatics and the lengths of bonds to hydrogen atoms are constrained using LINCS56 throughout all simulations. Before performing MD simulations for analysis, energy minimization is conducted for all initial structures, followed by equilibrium runs under NVT and NPT for 500 ps and 1 ns respectively.
To account for the induced-fit effect, Induced-Fit Docking (IFD) from the Schrödinger suite is the primary docking method used in this work. For comparison, we also tested the Glide XP docking from the Schrödinger suite and DiffDock to dock the two known type II inhibitors (Imatinib and Ponatinib) against DDR1 AF2 structure and the 15 classical DFG-out structures in the DDR1 rMSA ensemble.
4.2 Generation of rMSA AF2 ensemble, AF2-cluster ensemble and AF2-template structures
In this work, Colabfold 2 was employed to generate all the AF2-based ensembles and structures. The multiple sequence alignments (MSAs) were produced using mmseq2. For rMSA AF2 ensembles, the MSA depth was reduced to either 16 or 32 for each kinase. For each depth, 128 random seeds were utilized, and each seed produced 5 structural models via Colabfold, resulting in a total of 1280 rMSA AF2 structures per kinase. The structure with the highest pLDDT score among all 1280 was identified as the AF2 structure (native structure). Subsequently, any unphysical structures with an RMSD greater than 7 Å from the AF2 structure were discarded, culminating in the final rMSA AF2 ensembles.
AF2-cluster for DDR1, Abl1 and SrcK were run with default setups as provided in the ColabFold notebook in the original paper of AF2-cluster8.
The AF2-based homology modeling protocol, AF2-template (tAF2) method is implemented using Colabfold. For a given query sequence, tAF2 structures are generated by Colabfold upon uploading desired template structure and deactivating the Evoformer module. For each template, 5 AF2-template models are generated, and the last 3 structures exhibiting lower pLDDT will be discarded. The AF2-template (tAF2) method is employed here to transfer structure sets between homologous systems. To generate tAF2 structures for Abl1 in the classical DFG-out state, each of the 15 classical DFG-out structures from the DDR1 rMSA AF2 ensemble is used as a template, resulting in 30 tAF2 Abl1 structures in total. For Src kinase, a single representative tAF2 structure is generated using the “holo-model” DDR1 structure as the template.
4.3 AF2RAVE protocol
4.3.1 Regular space clustering on the rMSA AF2 ensemble
We used the same 14 collective variables (CVs) for regular space clustering as in the previous AF2RAVE work on kinases15. These CVs are pariwise distances, selected to describe the kinase conformations around the ATP-binding pocket and the A-loop.
4.3.2 Unbiased MD and SPIB
50 ns unbiased MD simulation was run for each AF2RAVE initial structure of DDR1 from Fig S2. The standard deviations of the 14 CVs are calculated after concatenating all the 12 unbiased MD trajectories. 8 CVs with standard deviations larger than 0.25 of the maximum standard deviation remain as the input features of the SPIB model. We conducted a parameter screening of the SPIB time lag, ranging from 1 ns to 40 ns with intervals of 1 ns. Eventually, we selected a time lag of 16 ns based on the performance of SPIB coordinates in representing physical features, including DFG-type and A-loop position.
4.3.3 PMF calculations from umbrella sampling
2D umbrella sampling is conducted along the two learnt SPIB coordinates, employing 11x11 windows. The bias potential equilibrium points are uniformly distributed in the SPIB latent space, with ranging from -0.8 to -0.1 and ranging from 0 to 0.8. The strength of the bias potential is set to 1000 . Each window originates from the structure closest to the window’s equilibrium point in Euclidean distance within the latent space among the 12 AF2RAVE initial structures and lasts for 100 ns (with the first 10 ns discarded in PMF calculation). For Abl1 kinase, we noticed a substantial proportion of the C helix breaking in the umbrella sampling trajectories (Fig S5). This finding aligns with earlier enhanced sampling investigations on DDR1 using metadynamics15, where the authors imposed restraints to prevent C helix breakage. In our study, we opted to exclude all umbrella sampling windows where the ratio of broken C helix exceeded 20% for the Abl1 PMF calculation (Fig S6). The WHAM algorithm is applied to bin and reweight the biased trajectories and compute the final PMF.
4.3.4 PMF calculations from unbiased MD
50 ns unbiased MD simulation was run starting from each structure in the 15 classical DFG-out structures in DDR1 rMSA AF2 ensemble. Upon discarding first 10 ns, all the unbiased trajectories are simply concatenated to calculate the Boltzmann distribution and PMF for each bin around the classical DFG-out basin in the latent space. For Abl1, unbiased simulations start from 30 AF2-template structures to calculate the PMF around the classical DFG-out basin.
4.3.5 Boltzmann ranks assignment for structures in AF2-based ensembles
After calculating the PMF value for each bin in the latent space, we projected AF2-generated structures into latent space. We then directly assign the PMF values of the corresponding bins to these AF2-generated structures.
We must acknowledge the limitations of the way we assigned PMF values to AF-generated candidate holo structures. First of all, the free energy profiles are derived from MD simulations, and the PMF values directly correspond to the MD structures. Here, we assumed that the latent space adequately represents the conformational changes of protein pocket within specific metastable states. Additionally, while the enrichment of holo structures in the top Boltzmann-ranked structures persists, the absolute PMF values and Boltzmann ranks of proper holo structures may fluctuate with the umbrella sampling setups, as depicted in Fig S7. Theoretically, the number of umbrella sampling windows and the simulation length should be sufficiently large for PMF convergence. However, there is always a trade-off between PMF accuracy and computational costs, so we opted to stick with the current setups.
4.4 Docking details
All the input structure for our docking experiments were first relaxed in solution with a MD energy minimization step. This work investigates two type I inhibitors (VX-680 and dasatinib) and two type II inhibitors (imatinib and ponatinib) by docking. For Glide XP docking or IFD, we used Ligprep in Maestro to prepare the ligand inputs from SMILES files. For DiffDock, the ligand inputs were directly provided as SMILES files.
4.4.1 Glide XP Docking
Glide XP docking experiments in this work were run with default setups in the Maestro software. Glide XP docking was performed for all four ligands on the AF2 structures of DDR1 or Abl1, as well as on the 15 classical DFG-out conformations of DDR1 in the rMSA AF2 ensemble.
4.4.2 Induced-Fit Docking (IFD)
For Induced-fit Docking, we used the Glide XP for initial docking, followed by Prime relaxation and final Glide XP docking. Parameters remain default in Maestro IFD. IFD was performed only for the type II ligands on the AF2 structures of DDR1 or Abl1, on the classical DFG-out conformations of DDR1 or Abl1 in the rMSA AF2 ensembles, as well as tAF2 structures of Abl1 and SrcK.
Given ponatinib’s backbone features, notably its lengthy and slender carbon-carbon triple bond, it exhibits reduced sensitivity to steric clashes, resulting in successful docking with the DDR1 “holo-model” structure at a ligand RMSD of 0.89 Å. Conversely, the “holo-model” structure struggles to accurately dock with imatinib using IFD (Fig S15B). We then employed an extended-sampling IFD approach by initially trimming the DFG-Phe residue from the “holo-model” structure. The trimmed residue is temporarily mutated to alanine during initial docking and later restored in subsequent Prime relaxation and final docking steps. Here, in the “holo-model” structure, we manually chose the DFG-Phe residue which is significantly hindered by holo-imatinib (Fig S16B). For generic systems lacking groundtruth information, broader screening of single residue trimming for the protein pocket may be necessary for this extended-sampling IFD method. In essence, it’s a trade-off between the quality of holo-like structure to dock with and the accuracy/complexity of the docking method.
4.4.3 Diffdock performace on DDR1 classical DFG-out in rMSA AF2 ensemble
[…]
@@ 4.4.3 Diffdock performace on DDR1 classical DFG-out in rMSA AF2 ensemble
Research in this publication was supported by the National Institute of General Medical Sciences of the National Institutes of Health under Award Number R35GM142719. The content is solely the responsibility of the authors and does not represent the official views of the National Institutes of Health. We thank UMD HPC’s Zaratan and NSF ACCESS (project CHE180027P) for computational resources. A.A. was supported by NCI-UMD Partnership for Integrative Cancer Research. P.T. is an investigator at the University of Maryland-Institute for Health Computing, which is supported by funding from Montgomery County, Maryland and The University of Maryland Strategic Partnership: MPowering the State, a formal collaboration between the University of Maryland, College Park and the University of Maryland, Baltimore. We thank UMD HPC’s Zaratan and NSF ACCESS (project CHE180027P) for computational resources. We thank Bodhi Vani, Dedi Wang, Zachary Smith and Anjali Verma for helpful discussions.
@@ 4.5 Notes
[…]